# Endoshare: A Source Available Solution to De-Identify and Manage Surgical Videos

### v1

@@ 7.3 CRediT Authorship Contribution Statement
Table 2: Exploratory factor analysis (EFA) of usability items. Bolded values ($\geq 0.40$) indicate the main factor for each item. Three factors emerged: (1) General Usability (usefulness and ease of use), (2) System Support & Error Management (recognition, recovery, help, consistency), and (3) Safety & Error Prevention (error prevention, control).
Domain Variable Factor 1 Factor 2 Factor 3
PU Speed 0.969 0.005 0.207
Performance 0.912 0.100 0.367
Productivity 0.946 -0.037 0.198
Effectiveness 0.901 0.164 0.362
Ease 0.899 0.206 0.365
Usefulness 0.995 0.005 0.053
PEOU Learning Ease 0.952 -0.119 0.076
Control 0.978 0.021 0.197
Clarity 0.978 0.070 -0.002
Flexibility 0.988 0.020 -0.023
Skill 0.988 0.020 -0.023
Ease of Use 0.977 0.020 0.087
Heur Visibility 0.280 0.498 0.503
Match 0.551 0.311 -0.246
Control 0.720 0.343 0.518
Consistency 0.058 0.672 0.091
Error Prevention 0.240 0.169 0.965
Recognition 0.280 0.943 -0.103
Flexibility 0.050 0.520 0.056
Error Recovery -0.191 0.969 0.037
Help -0.231 0.849 0.384
Table 3: Performance evaluation of EndoShare. The table summarizes the results of the technical performance assessment across three computing environments (Windows desktop, Linux workstation, and Mac laptop) and two processing modes (Advanced and Fast). (A) Descriptive statistics (mean ± SD) of processing times for videos of increasing duration (1, 30, and 60 minutes). (B) Linear mixed-effects model fitted on log-transformed times, where the Intercept represents the baseline condition (Advanced Mode, 1-minute video). The model showed significant main effects of processing mode ($p=0.001$) and video duration ($p<0.001$), with no significant interactions ($p>0.2$). (C) Paired comparisons indicated that Advanced Mode required 3.8× to 7.6× longer processing than Fast Mode, corresponding to an overall 6.2× (+519 %) increase, which did not reach statistical significance ($p=0.053$). (D) The Gamma generalized linear model (GLM) including machine type (the Intercept represents the baseline condition as Advanced Mode, 1-minute video, on the average desktop) confirmed the same pattern of effects as eTable 3B and identified a modest but significant difference between machines ($\chi^{2}(2)=6.37$, $p=0.041$), with the Mac laptop being the slowest and the Linux workstation the fastest.
(A) Summary Statistics
Machine Mode Video Mean $\pm$ SD (s)
Average Desktop (Windows) Advanced 1 min 39.4 $\pm$ 0.6
30 min 1202.7 $\pm$ 10.8
60 min 2417.1 $\pm$ 47.2
Fast 1 min 12.4 $\pm$ 0.2
30 min 290.8 $\pm$ 2.2
60 min 607.3 $\pm$ 6.6
High-end Desktop (Linux) Advanced 1 min 30.3 $\pm$ 2.2
30 min 1010.2 $\pm$ 10.7
60 min 1731.9 $\pm$ 66.6
Fast 1 min 7.5 $\pm$ 0.2
30 min 250.4 $\pm$ 25.8
60 min 437.7 $\pm$ 55.2
Laptop (macOS) Advanced 1 min 29.2 $\pm$ 2.1
30 min 6147.8 $\pm$ 2957.0
60 min 7718.7 $\pm$ 3630.5
Fast 1 min 6.6 $\pm$ 0.5
30 min 233.9 $\pm$ 66.3
60 min 697.2 $\pm$ 390.0
(B) Linear Mixed-Effects Model (log-time)
Fixed Effect Estimate ($\beta$) SE z p 95% CI
Intercept 3.486 0.317 10.980 <<0.001 [2.864, 4.108]
Mode = Fast -1.344 0.415 -3.236 0.001 [–2.158, –0.530]
Video = 30 min 4.092 0.415 9.854 <<0.001 [3.278, 4.906]
Video = 60 min 4.580 0.415 11.029 <<0.001 [3.766, 5.394]
[…]
(C) Paired comparisons (Advanced vs Fast) by video
Video length n GMR (Adv/Fast) % Difference t p-value
1 min 3 3.833997 +283.399743% 0.005218
30 min 3 7.597534 +659.753417% 0.082277
60 min 3 5.586656 +458.665581% 0.037322
Overall 3 6.192000 +519.200000% 0.052800
Bootstrap 95% CI (Overall GMR): [3.985, 14.819]
(D) Gamma GLM (log link) on Time
Fixed Effect Estimate ($\beta$) SE z p 95% CI
Intercept 3.510 0.336 10.432 <<0.001 [2.850, 4.169]
Machine = High-end Desktop (Linux) -0.302 0.291 -1.035 0.301 [–0.873, 0.269]
Machine = Laptop (MacOS) 0.380 0.291 1.304 0.192 [–0.191, 0.951]
Mode = Fast -1.309 0.412 -3.176 0.001 [–2.116, –0.501]
Video = 30 min 4.212 0.412 10.223 <<0.001 [3.405, 5.020]
Video = 60 min 4.606 0.412 11.178 <<0.001 [3.798, 5.413]
[…]
Likelihood-ratio test for Machine: $\chi^{2}(2)=6.374$, $p=0.041$.

### v2

@@ 7.3 CRediT Authorship Contribution Statement
Table 2: Performance evaluation of Endoshare. Panel (A) reports descriptive statistics for processing times, including geometric means (seconds) and slope estimates (seconds per minute, 95% CI) obtained from linear models through the origin fitted separately for each Machine $\times$ Mode. Panel (B) summarizes relative performance between fast and advanced modes using geometric mean ratios (GMR, fast / advanced) with bootstrap 95% confidence intervals and slope ratios. Consistent slope and GMR values across video lengths indicate proportional (linear) scaling of processing time with video duration and a stable relative speed advantage of fast mode across hardware configurations.
Machine / Mode / Video Geometric Mean (s) Slope (s/min) [95% CI]
Average Desktop (Windows) Advanced 1 min 39.4 40.25 (39.63–40.86)
30 min 1202.7
60 min 2416.8
Fast 1 min 12.4 10.04 (9.96–10.13)
30 min 290.8
60 min 607.3
High-end Desktop (Linux) Advanced 1 min 30.2 29.82 (28.96–30.65)
30 min 1010.2
60 min 1731.0
Fast 1 min 7.5 7.48 (6.77–8.18)
30 min 249.6
60 min 435.3
Laptop (macOS) Advanced 1 min 29.1 137.60 (101.98–190.23)
30 min 5729.8
60 min 7213.7
Fast 1 min 6.6 10.14 (5.93–16.02)
30 min 227.3
60 min 619.2
(A) Fast vs Advanced
Machine GMR (Fast/Advanced) [95% CI] Slope ratio (Fast/Advanced)
Average Desktop (Windows) 0.269 (0.265–0.274) 0.249
High-end Desktop (Linux) 0.250 (0.225–0.274) 0.251
Laptop (macOS) 0.120 (0.092–0.158) 0.074